# Gaga: Group Any Gaussians via 3D-aware Memory Bank

### v2

@@ Abstract
Figure 1: Gaga groups any Gaussians in an open-world 3D scene and renders multi-view consistent segmentation (pixels of the same region across views are represented with the same color). By employing a 3D-aware memory bank, we eliminate the label inconsistency that exists in 2D segmentation predicted by foundational models and assign each mask across different views a universal group ID. This enables the process of lifting 2D segmentation to a consistent 3D segmentation. Gaga produces accurate 3D object segmentation, achieving high-quality results for downstream applications such as scene manipulation (e.g. changing the cushion’s color of the footstool to maroon).
@@ 1 Introduction
Figure 2: Comparison of rendered segmentation. Contrastive learning-based methods, such as OmniSeg [33], do not provide unique mask labels to each segmentation group, leading to inconsistencies across multiple views (e.g., the coffee table). Gaussian Grouping [32] addresses multi-view segmentation by utilizing a video tracker, but it often misidentifies objects when similar items are present (e.g., the leather sofa) and struggles with significant camera perspective changes. In contrast, Gaga ensures multi-view consistent segmentation masks, overcoming these limitations.
@@
Effective open-world 3D segmentation is essential for scene understanding and manipulation. Despite notable advancements in 2D open-world segmentation techniques, exemplified by Segment Anything (SAM) [16] and EntitySeg [23], extending these methodologies to the realm of 3D encounters the challenge of ensuring consistent mask label assignment across multi-view images. Specifically, masks of the same object across different views may have different mask label IDs, as the 2D class-agnostic segmentation model independently processes the multi-view images. Naively lifting these inconsistent 2D masks to 3D introduces ambiguity and leads to inferior results in 3D scene segmentation.
@@
Existing works [33, 15, 32, 11, 9] address this issue either by contrastive learning or video object tracking. One line of works build upon 3D scenes represented as 3D Gaussians [13]. In this approach, a feature vector is learned for each 3D Gaussian through a contrastive learning paradigm. The method encourages similar features for 3D Gaussians whose projections fall within the same segmentation mask, while pushing apart those belonging to different segmentation regions. Segmentation maps for each view are then obtained by clustering these feature vectors. However, these methods do not assign a specific mask label to each 3D segmentation group, resulting in inconsistent multi-view segmentation, as shown in Fig. 2. Alternatively, some methods [32, 11] tackle multi-view inconsistency in 3D scene segmentation by treating multi-view images as a video sequence and adopting an off-the-shelf video object tracking method [7]. Nevertheless, this design assumes minimal view changes between multi-view images—a condition that may not always hold in real-world 3D scenes, particularly when input views are sparse. Consequently, these methods often struggle with similar objects or occluded objects that intermittently disappear and reappear in the sequence, as shown in Fig. 2.
@@
In this paper, we identify a fundamental limitation in existing open-world 3D scene segmentation methods that leads to 3D inconsistency: the inadequate exploitation of 3D information inherently provided by the scene. Our main intuition is that masks of the same object across different views shall correspond to the same group of 3D Gaussians. Hence, we can assign identical universal mask IDs to masks from different views when there is a large overlap between their corresponding 3D Gaussian groups.
@@
Based on this intuition, we propose Gaga, which groups any 3D Gaussians and renders consistent 3D class-agnostic segmentation across different views. Given a collection of posed RGB images, we first employ Gaussian Splatting [13] to reconstruct the 3D scene and extract 2D masks using an open-world segmentation method [16, 23]. Subsequently, we iteratively build a 3D-aware memory bank that collects and stores the Gaussians grouped by category. For each input view, we project each 2D mask into 3D space using camera parameters and search the memory bank for the category with the largest overlapping with the unprojected mask. Based on the degree of overlapping, we either assign the mask to an existing category or create a new one. Finally, following the mask association process described above, we can get a set of multi-view consistent 2D segmentation masks. Finally, we learn a feature vector (i.e., identity encoding [32]) for each 3D Gaussian that encodes its category information. Specifically, we splat the feature vectors onto 2D image and decode a segmentation map through a linear layer. The predicted masks are then compared with the segmentation masks obtained from our 3D-aware memory bank for supervision. This approach ensures that our identity encoding remains multi-view consistent, thanks to the consistent segmentation masks across different viewpoints. Our contributions include:
  • We propose a framework that reconstructs and segments 3D scenes using inconsistent 2D masks generated by open-world segmentation models.
  • To resolve the inconsistency of 2D masks across views, we design a 3D-aware memory bank that collects Gaussians of the same semantic group. This memory bank is then employed to align 2D masks across diverse views.
  • We show that the proposed method can effectively leverage any 2D class-agnostic segmentation masks, making it readily applicable for synthesizing novel view images and segmentation masks.
  • Comprehensive experiments on diverse datasets and challenging scenarios, including sparse input views, demonstrate the effectiveness of the proposed method both qualitatively and quantitatively.
@@ 2 Related Work
Segment and Tracking Anything in 2D. Segment Anything (SAM) [16] and EntitySeg [23] demonstrate the effectiveness of large-scale training in image segmentation, thus establishing a pivotal foundation for open-world segmentation methods. Subsequent studies [31, 8, 7] further extend the applicability of SAM to video data by leveraging video object segmentation algorithms to propagate SAM masks. Conversely, acquiring data for training their 3D counterparts poses a challenge, given that existing large-scale 3D datasets with annotated segmentation [27, 10] primarily focus on indoor scenarios.
@@
NeRF-based 3D Segmentation. Neural Radiance Fields (NeRFs) [21] model scenes as continuous volumetric functions, learned through neural networks that map 3D coordinates to scene radiance. This approach facilitates the capture of intricate geometric details and the generation of photorealistic renderings, offering novel view synthesis capabilities. Semantic-NeRF [35] initiates the incorporation of semantic information into NeRFs and enables the generation of semantic masks for novel views. Note that semantic segmentation masks do not face the challenge of ambiguous mask ID across views. Numerous methods expand the scope by introducing instance modeling and matching instance masks relying on existing 3D bounding boxes [20, 12], resorting to the cost-based linear assignment during training [26, 29] or directly training instance-specific MLPs [17]. However, most of these methods are developed based on ground truth segmentation and tailored for scene modeling within specific domains. They often entail high computational costs and lack substantial evidence of their performance in open-world scenarios.
Leveraging SAM’s open-world segmentation capability, SA3D [3] endeavors to recover a 3D consistent mask by tracing 2D masks across adjacent views with user guidance. Similarly, Chen et al. [6] distill SAM encoder features into 3D and query the decoder. In contrast, Gaga achieves multi-view consistency without user intervention, offering segmentation for all objects rather than an instance. Garfield [15] densely samples SAM masks and trains a scale-conditioned affinity field supervised on the scale of each mask deprojected to 3D. However, this work focuses on 3D clustering rather than segmentation.
Gaussian-based 3D Segmentation. As an alternative to NeRF and its variants [21, 4, 22, 28], Gaussian Splatting [13, 5, 30, 34] has recently emerged as a powerful approach to reconstruct 3D scenes via real-time radiance field rendering. By representing the scene as 3D Gaussians from posed images, it achieves photorealistic novel view synthesis with high reconstruction quality and efficiency. Additionally, manipulating 3D Gaussians for scene editing is more straightforward compared to NeRF’s representation.
@@
SAGA [2] renders a 2D SAM feature map and uses a SAM guidance loss to learn 3D segmentation from the ambiguous 2D masks. Similar to [3], this method requires user input and only provides segmentation for one object at a time. Feature 3DGS [36] distills LSeg [18] and SAM features to 3D Gaussians and decodes rendered features to obtain segmentation. However, it fails to provide consistent segmentation across views. Gaussian Grouping [32] and CoSSegGaussians [11] use a video object tracker [7] to associate masks across different views. However, in scenarios with significant changes in camera poses between frames, such approaches struggle to maintain accuracy. OmniSeg3D [33] and Click-Gaussians [9] use contrastive learning-based method to attach each Gaussian with a feature, while the features of two Gaussians are encouraged to be similar if their projections are within the same segmentation mask. However, such methods do not provide segmentation mask labels for each segmented Gaussian group.
@@ 3.1 Preliminaries
Gaussian Splatting. Gaussian Splatting [13] has significantly advanced the 3D representation field by combining the benefits of implicit and explicit 3D representations. Specifically, a 3D scene is parameterized by a set of 3D Gaussians $\{G_{i}\}$. Each Gaussian $G_{i}=\{p_{i},s_{i},q_{i},\alpha_{i},c_{i}\}$ is defined by its position $p_{i}=\{x,y,z\}\in\mathbb{R}^{3}$, scale $s_{i}\in\mathbb{R}^{3}$, orientation $q\in\mathbb{R}^{4}$, opacity $\alpha\in\mathbb{R}$ and color features $c$ encoded by spherical harmonics (SH) coefficients.
Gaussian Splatting employs the splatting pipeline, wherein 3D Gaussians are projected onto the 2D image space using the world-to-frame transformation matrix corresponding to each camera pose. Gaussians projected to the same coordinates $(x,y)$ (represented as $i\in N$) are blended in depth order and weighted by their opacity $\alpha$ to produce the color $c_{x,y}$ of each pixel:
$c_{x,y}=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}).$ (1)
Identity Encoding. Identity encoding ($e_{i}$[32] aims to assign a universal label to each 3D Gaussian for 3D scene segmentation. An identity encoding is a 16-dimensional feature vector attached to a 3D Gaussian. It can be decoded to a segmentation mask ID through a combination of linear and SoftMax layers. The segmentation label of a pixel (denoted as $m_{x,y}$) can then be computed by:
The learning process for identity encoding involves optimizing the feature vectors by comparing the rendered masks with the ground truth 2D segmentation masks.
3.2 Grouping via 3D-aware Memory Bank
Given a set of posed images, we aim to reconstruct a 3D scene with semantic labels for segmentation rendering. To this end, we first leverage Gaussian Splatting for scene reconstruction. We then employ open-vocabulary 2D segmentation methods such as SAM [16] or EntitySeg [23] to predict class-agnostic segmentation for each input view.
However, because the segmentation model processes each input view independently, the resulting masks are not naturally multi-view consistent. To resolve this issue, existing works [32, 11] carry out a mask association process that tries to consolidate inconsistent segmentation masks from different views. For example, GaussianGrouping [32] and CoSSegGaussians [11] apply a video tracker to associate inconsistent 2D masks of different views, assuming similarity between nearby input views. This assumption, however, may not hold for all 3D scenes, particularly when the input views are sparse, as demonstrated in Fig. 2.
Gaga is inspired by the fundamental disparity between the task of mask association across multiple views and object tracking in a video: the latter does not exploit 3D information in the scene. To reliably generate consistent masks across different views, we propose a method that leverages 3D information without relying on any assumptions about the input views. Our key insight is that masks belonging to the same object in different views shall correspond to the same Gaussians in the 3D space. Consequently, these Gaussians should be grouped together and assigned with an identical group ID.
@@ 3.2 Grouping via 3D-aware Memory Bank
$Q_{1}=\text{Quantile}(D,0.25),\quad Q_{3}=\text{Quantile}(D,0.75).$ (3)
@@
Note that we do not use the IoU of $\mathcal{G}(m)$ and $\mathcal{G}_{i}$ for the following reason: as more frames are processed, more Gaussians will be added to the memory bank, which means that $\mathcal{G}_{i}$ will become larger. Consequently, the IoU threshold needs to be adjusted frequently. In contrast, our overlapping formulation in Eq. 5 is independent of Gaussian number in the memory bank, avoiding manual threshold adjustment.
@@
After the group ID assigning process, masks projected by the same group of Gaussians are supposed to have the same group ID across different views. We splat the group IDs of 3D Gaussians in the memory bank to obtain pseudo ground truth segmentation masks, which are used to learn a 16-dimension feature vector (i.e., identity encoding [32]) for each Gaussian. Specifically, we initialize the identity encoding to all zeros for each 3D Gaussian. At each training iteration, we splat the identity encoding given a camera pose to formulate a 2D feature map. A linear layer followed by a SoftMax function are employed to predict a 2D segmentation map given the rendered feature map. Finally, we compute the Cross Entropy loss between the predicted segmentation maps and those obtained by rendering Gaussians in the memory bank, as discussed above.
@@ 4.1 Experimental Setup
Datasets. We experiment with various datasets across diverse scenarios to demonstrate the performance of Gaga. For quantitative comparison, we use a scene understanding dataset LERF-Mask [32], along with two indoor scene datasets: Replica [27] and ScanNet [10]. Additionally, we showcase the robustness of Gaga against variations in training image quantity by sparsely sampling the Replica dataset. We present visual comparison results on the commonly used scene reconstruction dataset, MipNeRF 360 [1]. The quantitative and qualitative evaluations are conducted on the test set, i.e., novel view synthesis results. Details about datasets can be found in the supplementary material.
Evaluation Metrics. Similarly to prior work [32], mIoU and boundary IoU (mBIoU) are used to evaluate the LERF-Mask dataset. To evaluate the multi-view consistency of 3D segmentation masks, we select eight scenes from the Replica dataset and seven scenes from the ScanNet dataset and use the ground-truth panoptic segmentation for evaluation, disregarding class information. The evaluation process is detailed in the supplementary material. Note that only Gaussian Grouping and Gaga provide 3D mask IDs to render multi-view consistent segmentation masks, we only compare these two methods. To handle differences between predicted and ground truth mask labels, we calculate the best linear assignment based on IoU. Moreover, with IoU = 0.5 as the criterion, we report precision and recall to further evaluate the accuracy of predicted masks. Please note that all numbers in the tables are expressed as percentages.
Implementation Details. We use SAM [16] and EntitySeg [23] with the Hornet-L backbone to obtain open-world 2D segmentation. We preprocess the generated raw masks following the method outlined in [23], prioritizing those with higher confidence scores by ranking them accordingly. Masks with confidence scores below 0.5 are discarded. For all experiments, we train vanilla Gaussian Splatting [13] for 30K iterations and train the identity encoding for 10K iterations with all other parameters frozen. We train baseline methods following their official guidance for 40K iterations for fair comparisons. We set the overlap threshold for declaring a new group ID as 0.1.

### v3

# Group Any Gaussians via 3D-aware Memory Bank

@@ Abstract
Figure 1: Gaga groups any Gaussians in an open-world 3D scene and renders multi-view consistent segmentation (pixels of the same region across views are represented with the same color). By employing a 3D-aware memory bank, we eliminate the label inconsistency that exists in 2D segmentation predicted by foundational models and assign each mask across different views a universal group ID. This enables the process of lifting 2D segmentation to a consistent 3D segmentation. Gaga produces accurate 3D object segmentation, achieving high-quality results for downstream applications such as scene manipulation (e.g. changing the cushion’s color of the footstool to maroon)
@@ 1 Introduction
Effective open-world 3D segmentation is essential for scene understanding and manipulation. Despite notable advancements in 2D open-world segmentation techniques, exemplified by Segment Anything (SAM) sam and EntitySeg entityseg , extending these methodologies to the realm of 3D encounters the challenge of ensuring consistent mask label assignment across multi-view images. Specifically, masks of the same object across different views may have different mask label IDs, as the 2D class-agnostic segmentation model independently processes the multi-view images. Naively lifting these inconsistent masks to 3D introduces ambiguity and leads to inferior results in 3D scene segmentation.
Figure 2: Comparison of rendered segmentation. Contrastive learning-based methods, such as OmniSeg omniseg3d , do not provide unique mask labels to each segmentation group, leading to inconsistencies across multiple views (e.g., the coffee table). Gaussian Grouping gg addresses multi-view segmentation by utilizing a video tracker, but it often misidentifies objects when similar items are present (e.g., the leather sofa) and struggles with significant camera perspective changes. In contrast, Gaga ensures multi-view consistent segmentation masks, overcoming these limitations.
@@
Existing works omniseg3d ; garfield ; gg ; cosseg ; clickgaussian address this issue either by contrastive learning or video object tracking. One line of work builds upon 3D scenes represented as 3D Gaussians 3dgs . In this approach, a feature vector is learned for each 3D Gaussian through a contrastive learning paradigm. The method encourages similar features for 3D Gaussians whose projections fall within the same segmentation mask, while pushing apart those belonging to different segmentation regions. The segmentation maps for each view are then obtained by clustering these feature vectors. However, these methods do not assign a specific mask label to each 3D segmentation group, resulting in inconsistent multi-view segmentation, as shown in Fig. 2. Alternatively, some methods gg ; cosseg tackle multi-view inconsistency in 3D scene segmentation by treating multi-view images as a video sequence and adopting an off-the-shelf video object tracking method deva . Nevertheless, this design assumes minimal view changes between multiview images, a condition that may not always hold in real-world 3D scenes, particularly when input views are sparse. Consequently, these methods often struggle with similar objects or occluded objects that intermittently disappear and reappear in the sequence, as shown in Fig. 2.
@@
We identify a fundamental limitation in existing open-world 3D scene segmentation methods that leads to 3D inconsistency: the inadequate exploitation of 3D information inherently provided by the scene. Our main intuition is that masks of the same object across different views shall correspond to the same group of 3D Gaussians. Hence, we can assign identical universal mask IDs to masks from different views when there is a large overlap between their corresponding 3D Gaussian groups.
@@
Based on this intuition, we propose Gaga, which groups any 3D Gaussians and renders consistent 3D class-agnostic segmentation across different views. Given a collection of posed RGB images, we first employ Gaussian Splatting 3dgs to reconstruct the 3D scene and extract 2D masks using an open-world segmentation method sam ; entityseg . Subsequently, we iteratively build a 3D-aware memory bank that collects and stores the Gaussians grouped by category. For each input view, we project each 2D mask into a 3D space using camera parameters and search the memory bank for the category with the largest overlap with the unprojected mask. Based on the degree of overlap, we assign the mask to an existing category or create a new one. Finally, following the mask association process described above, we can get a set of multi-view consistent 2D segmentation masks. Finally, we learn a feature vector (i.e., identity encoding gg ) for each 3D Gaussian that encodes its category information. Specifically, we splat the feature vectors onto the 2D image and decode a segmentation map through a linear layer. The predicted masks are then compared with the segmentation masks obtained from our 3D-aware memory bank for supervision. This approach ensures that our identity encoding remains multi-view consistent, thanks to the consistent segmentation masks across different viewpoints.
The contributions of this work are: 1). We propose a framework for reconstructing and segmenting 3D scenes from inconsistent 2D masks generated by open-world segmentation models. 2). To resolve the inconsistency of 2D masks across views, we design a 3D-aware memory bank that collects Gaussians of the same semantic group. This memory bank is then used to align 2D masks across diverse views. 3). We show that the proposed method can effectively leverage any 2D class-agnostic segmentation mask, making it readily applicable for synthesizing novel view images and segmentation masks. 4.) Experiments on diverse datasets and challenging scenarios, including sparse input views, demonstrate the effectiveness of the proposed method qualitatively and quantitatively.
@@ 2 Related Work
Segment and Tracking Anything in 2D. Segment Anything (SAM) sam and EntitySeg entityseg demonstrate the effectiveness of large-scale training in image segmentation, thus establishing a pivotal foundation for open-world segmentation methods. Subsequent studies track ; sata ; deva further extend the applicability of SAM to video data by leveraging video object segmentation algorithms to propagate SAM masks. Conversely, acquiring data for training their 3D counterparts poses a challenge, given that existing large-scale 3D datasets with annotated segmentation replica ; scannet focus on indoor scenarios.
@@
NeRF-based 3D Segmentation. Neural Radiance Fields (NeRFs) nerf model scenes as continuous volumetric functions, learned through neural networks that map 3D coordinates to scene radiance. This approach facilitates the capture of intricate geometric details and the generation of photorealistic renderings, offering novel view synthesis capabilities. Semantic-NeRF semantic_nerf initiates the incorporation of semantic information into NeRFs and enables the generation of semantic masks for novel views. Note that semantic segmentation masks do not face the challenge of ambiguous mask ID across views. Numerous methods expand the scope by introducing instance modeling and matching instance masks relying on existing 3D bounding boxes instance_nerf ; panoptic_nerf , resorting to the cost-based linear assignment during training panoptic_lifiting ; dm_nerf or directly training instance-specific MLPs pnf . However, most of these methods are developed based on ground truth segmentation and tailored for scene modeling within specific domains. They often entail high computational costs and lack substantial evidence of their performance in open-world scenarios. Leveraging SAM’s open-world segmentation capability, SA3D sa3d endeavors to recover a 3D consistent mask by tracing 2D masks across adjacent views with user guidance. Similarly, Chen et al. interactive distill SAM encoder features into 3D and query the decoder. In contrast, Gaga achieves multi-view consistency without user intervention, offering segmentation for all objects rather than an instance. Garfield garfield densely samples SAM masks and trains a scale-conditioned affinity field supervised on the scale of each mask deprojected to 3D, which focuses on 3D clustering rather than segmentation.
@@
Gaussian-based 3D Segmentation. As an alternative to NeRF and its variants nerf ; tensorf ; instant ; nerfstudio , Gaussian Splatting 3dgs ; chen2024survey ; recent ; Yu2023MipSplatting has recently emerged as a powerful approach to reconstruct 3D scenes via real-time radiance field rendering. By representing the scene as 3D Gaussians from posed images, it achieves photorealistic novel view synthesis with high reconstruction quality and efficiency. Additionally, manipulating 3D Gaussians for scene editing is more straightforward compared to NeRF’s representation. SAGA saga renders a 2D SAM feature map and uses a SAM guidance loss to learn 3D segmentation from the ambiguous 2D masks. Similar to sa3d , this method requires user input and only provides segmentation for one object at a time. Feature 3DGS feature3dgs distills LSeg lseg and SAM features to 3D Gaussians and decodes rendered features to obtain segmentation. However, it fails to provide consistent segmentation across views. Gaussian Grouping gg and CoSSegGaussians cosseg use a video object tracker deva to associate masks across different views. However, in scenarios with significant changes in camera poses between frames, such approaches struggle to maintain accuracy. OmniSeg3D omniseg3d and Click-Gaussians clickgaussian use contrastive learning-based method to attach each Gaussian with a feature, while the features of two Gaussians are encouraged to be similar if their projections are within the same segmentation mask. However, such methods do not provide segmentation mask labels for each segmented Gaussian group.
@@ 3.1 Preliminaries
Gaussian Splatting. Gaussian Splatting 3dgs has significantly advanced the 3D representation field by combining the benefits of implicit and explicit 3D representations. Specifically, a 3D scene is parameterized by a set of 3D Gaussians $\{G_{i}\}$. Each Gaussian $G_{i}=\{p_{i},s_{i},q_{i},\alpha_{i},c_{i}\}$ is defined by its position $p_{i}=\{x,y,z\}\in\mathbb{R}^{3}$, scale $s_{i}\in\mathbb{R}^{3}$, orientation $q\in\mathbb{R}^{4}$, opacity $\alpha\in\mathbb{R}$ and color features $c$ encoded by spherical harmonics (SH) coefficients. Gaussian Splatting employs the splatting pipeline, wherein 3D Gaussians are projected onto the 2D image space using the world-to-frame transformation matrix corresponding to each camera pose. Gaussians projected to the same coordinates $(x,y)$ (represented as $i\in N$) are blended in depth order and weighted by their opacity $\alpha$ to produce the color $c_{x,y}$ of each pixel: $c_{x,y}=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$.
3.2 Grouping Any Gaussians via 3D-aware Memory Bank
Given a set of posed images, we aim to reconstruct a 3D scene with semantic labels for segmentation rendering. To this end, we first leverage Gaussian Splatting for scene reconstruction. We then employ open-vocabulary 2D segmentation methods such as SAM sam or EntitySeg entityseg to predict class-agnostic segmentation for each input view. However, because the segmentation model processes each input view independently, the resulting masks are not naturally multi-view consistent. To resolve this issue, existing works gg ; cosseg carry out a mask association process that tries to consolidate inconsistent segmentation masks from different views. For example, GaussianGrouping gg and CoSSegGaussians cosseg apply a video tracker to associate inconsistent 2D masks of different views, assuming similarity between nearby input views. This assumption, however, may not hold for all 3D scenes, particularly when the input views are sparse, as demonstrated in Fig. 2.
Gaga is inspired by the fundamental disparity between the task of mask association across multiple views and object tracking in a video: the latter does not exploit 3D information in the scene. To reliably generate consistent masks across different views, we propose a method that leverages 3D information without relying on any assumptions about the input views. Our key insight is that masks belonging to the same object in different views shall correspond to the same Gaussians in the 3D space. Consequently, these Gaussians should be grouped together and assigned an identical group ID.
@@ 3.2 Grouping Any Gaussians via 3D-aware Memory Bank
$Q_{1}=\text{Quantile}(D,0.25),\quad Q_{3}=\text{Quantile}.(D,0.75).$ (1)
@@
Note that we do not use the IoU of $\mathcal{G}(m)$ and $\mathcal{G}_{i}$ for the following reason: as more frames are processed, more Gaussians will be added to the memory bank, which means that $\mathcal{G}_{i}$ will become larger. Consequently, the IoU threshold needs to be adjusted frequently. In contrast, our overlapping formulation in Eq. 3 is independent of Gaussian number in the memory bank, avoiding manual threshold adjustment. In practice, We set the overlap threshold for declaring a new group ID as 0.1.
@@
After the group ID assigning process, masks projected by the same group of Gaussians are supposed to have the same group ID across different views. We splat the group IDs of 3D Gaussians in the memory bank to obtain pseudo ground truth segmentation masks, which are used to learn a 16-dimensional feature vector (i.e., identity encoding gg ) for each Gaussian. Identity encoding ($e_{i}$gg aims to assign a universal label to each 3D Gaussian for 3D scene segmentation. It can be decoded to a segmentation mask ID through a combination of linear and SoftMax layers. The segmentation label of a pixel (denoted as $m_{x,y}$) is computed by:
The learning process for identity encoding involves optimizing the feature vectors by comparing the rendered masks with the ground truth 2D segmentation masks. Specifically, we initialize the identity encoding to all zeros for each 3D Gaussian. At each training iteration, we splat the identity encoding given a camera pose to formulate a 2D feature map. A linear layer followed by a SoftMax function are employed to predict a 2D segmentation map given the rendered feature map. Finally, we compute the Cross Entropy loss between the predicted segmentation maps and those obtained by rendering Gaussians in the memory bank, as discussed above.
@@ 4.1 Experimental Setup
Datasets. For quantitative comparison, we use a scene understanding dataset LERF-Mask gg , along with two indoor scene datasets: Replica replica and ScanNet scannet . Additionally, we showcase the robustness of Gaga against variations in training image quantity by sparsely sampling the Replica dataset. We present visual comparison results on the commonly used scene reconstruction dataset, MipNeRF 360 mipnerf . The quantitative and qualitative evaluations are conducted on the test set, i.e., novel view synthesis results. Details about datasets can be found in the supplementary material.
Evaluation Metrics. Similarly to prior work gg , mIoU and boundary IoU (mBIoU) are used to evaluate the LERF-Mask dataset. To evaluate the multi-view consistency of 3D segmentation masks, we select eight scenes from the Replica dataset and seven scenes from the ScanNet dataset and use the ground-truth panoptic segmentation for evaluation, disregarding class information. The evaluation process is detailed in the supplementary material. Note that only Gaussian Grouping and Gaga provide 3D mask IDs to render multi-view consistent segmentation masks, we only compare these two methods. To handle differences between predicted and ground truth mask labels, we calculate the best linear assignment based on IoU. Moreover, with IoU = 0.5 as the criterion, we report precision and recall to further evaluate the accuracy of predicted masks. Please note that all numbers in the tables are expressed as percentages.
Implementation Details. We use SAM sam and EntitySeg entityseg with the Hornet-L backbone to obtain open-world 2D segmentation. We preprocess the generated raw masks following the method outlined in entityseg , prioritizing those with higher confidence scores by ranking them accordingly. Masks with confidence scores below 0.5 are discarded. For all experiments, we train vanilla Gaussian Splatting 3dgs for 30K iterations and train the identity encoding for 10K iterations with all other parameters frozen. We train baseline methods following their official guidance for 40K iterations for fair comparisons.
4.2 Open-vocabulary 3D Query on LERF-Mask
Table 1: 3D query on LERF-Mask. Gaga outperforms previous approaches, showcasing favorable performance in mIoU and BIoU with both segmentation models.